Circular No. 6653

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUNGRAZING COMETS Further to IAUC 6650, C. St. Cyr, NASA Goddard Space Flight Cen-

ter, has also reported detections of other comets in images obtained with the SOHO coronagraphs, the following approximate discovery positions being summarized from the more complete reductions from C2 and C3 data by G. V. Williams on MPEC 1997-J06 to 1997-J09:

1996 UT  α_{2000}  δ_{2000}
Aug. 20.933  9^{h}34.^{m}9  +10^{o}17′  C/1996 Q2
30.073  10 14.2  + 7 49  C/1996 Q3
Sept. 22.265  11 28.5  − 1 32  C/1996 S3
Dec. 22.490  18 12.9  −26 21  C/1996 Y1

C/1996 Q2 and C/1996 Y1 were discovered by Shane Stezelberger, SOHO-LASCO operations, C/1996 Q3 and C/1996 S3 by D. Biesecker, University of Birmingham. Orbital computations by the undersigned on the same MPEC s suggest that all these SOHO comets are Kreutz sungrazers, with T = Aug. 22.13, Aug. 30.87, Sept. 23.58 and Dec. 23.26, respectively.

COMET C/1996 J1 (EVANS-DRINKWATER) S. Nakano, Sumoto, Japan, reports that low-altitude CCD observations

by J. Kobayashi (using a 0.41-m reflector at Kumamoto Civil Astronomical Observatory) on May 5.8 UT show a bright condensation (“nucleus A”) of m_{1} = 9.8 and a fainter condensation (“nucleus B”) of m_{1} = 12.9, separated from the former by ∼ 81′′ in p.a. 72^{o}. There is a tail 5′ long in p.a. 260^{o}, probably from nucleus A and thus influencing its brightness estimate above. The comet’s splitting was confirmed by A. Sugie (using a 0.60-m reflector at Dynic Astronomical Observatory) on May 9.8, with the condensations at m_{1} = 12.6 and 13.7, respectively. From its better fit with the 1996 observations, Nakano suggests that nucleus B is the principal one, a point that is consistent with its earlier perihelion time (by 0.12 day if one represents the observations with otherwise identical orbital elements).

Ephemeris from the orbital elements (nucleus B) on MPEC 1997-J10:

1997TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
May 2 0^{h}55.^{m}73 +23^{o}10.^{′}5 3.047 2.177 25.^{o}1 11.^{o}3 12.3
12 1 14.84 +25 25.6 3.113 2.282 28.9 12.4 12.5
22 1 32.83 +27 28.9 3.170 2.388 33.2 13.4 12.8
June 1 1 49.70 +29 21.5 3.215 2.495 38.0 14.5 13.0
11 2 05.39 +31 04.4 3.249 2.602 43.1 15.5 13.2
1997 May 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden